(1.1) Let U and V be the planes given by: U : λx + 5y − 2λz − 3 = 0,

V : −λx + y + 2z + 1 = 0.

Determine for which value(s) of λ the planes U and V are:

(a) orthogonal,

(b) Parallel.

(1.2) Find an equation for the plane that passes through the origin (0, 0, 0) and is parallel to the (3) plane −x + 3y − 2z = 6.

(1.3) Find the distance between the point (−1, −2, 0) and the plane 3x − y + 4z = −2.


1
Expert's answer
2021-07-14T11:28:06-0400

1.

a)

"-\lambda^2+5-4\lambda=0"

"\lambda_1=1, \lambda_2=4"


b)

The planes are parallel when

"\lambda/-\lambda=5=-2\lambda/2"

So. such "\lambda" does not exist.


2.

−x + 3y − 2z =0


3.

"d=\frac{|-3+2+2|}{\sqrt{3^2+1+4^2}}=1/\sqrt{26}"